What is the equation of a line that passes through point (1, -3) and has a slope of -2? The answer needs to be written in slope-intercept form.
Answer by mananth(16946)

What is the equation of a line that passes through point (1, -3) and has a slope of -2? The answer needs to be written in slope-intercept form.
..
find b in the equation y = mx+b
plug values of x,y & m
-3=-2*1+b
-3=-2+b
b=-3+2
b=-1
...
y= -2x-1
..
m.ananth@hotmail.ca